OpenStudy (anonymous):

Half-Angle Identities: How would i solve sin(13pi/12)? Don't know what to do with the 13, since it is not part of the Unit Circle.

OpenStudy (anonymous):

Consider that 13 = 9 + 4

OpenStudy (anonymous):

9pi/12 and 4pi/12 should be on the unit circle.

OpenStudy (anonymous):

you can use the identity: \[ \sin(x+y) = \sin(x)\cos(y) + \sin(y)\cos(x) \]

OpenStudy (anonymous):

i have to do that with in the equation? the sinx/2

OpenStudy (anonymous):

Okay then try the halfangle formula.